OpenStudy (anonymous):

Simplify the given expression: 4 divided by the quantity of 3 minus 2i

zepdrix (zepdrix):

Multiply by the `conjugate` of the denominator, \[\Large\rm \frac{4}{3-2\mathcal i}\color{royalblue}{\left(\frac{3+2\mathcal i}{3+2\mathcal i}\right)}\]

zepdrix (zepdrix):

multiplying complex conjugates gives us the `sum of their squares`, \[\Large\rm (a-\mathcal i b)(a+\mathcal i b)=a^2+b^2\]So that's what happening in the denominator,\[\Large\rm \frac{4(3+2\mathcal i)}{3^2+2^2}\]

zepdrix (zepdrix):

Just requires a little bit of work to simplify it from there.